OpenStudy (anonymous):

1/(11^20) - 1/(11^21) = 10/(11^21) Can someone tell me the steps taken to get to that answer?

OpenStudy (anonymous):

Make the denominators the same, and it should make sense.

OpenStudy (anonymous):

1/(11^20) - 1/(11^21) = 10/(11^21) \(\large \frac{1}{11^{20}}-\frac{1}{11^{21}} \) \(\large \frac{1}{11^{20}}\color {red}{\cdot \frac{11}{11}}-\frac{1}{11^{21}} \)

OpenStudy (anonymous):

now the denominators are the same, you all you need to do is subtract the numerators.

OpenStudy (anonymous):

Whoa. Coloured text. How'd you do that?

OpenStudy (anonymous):

I've been wanting to do that all day.

OpenStudy (anonymous):

right click "show math" ---> "as tex" insert this " \color {red}" into your LaTex....

OpenStudy (anonymous):

Ahhhh...fantastic. Thank you.

OpenStudy (anonymous):

yw... :)

OpenStudy (anonymous):

hmmm.. i was gonna point you to the Latex group here on OS but apparently it's gone... if you wanna "see" the code for Latex , see this: http://www.codecogs.com/latex/eqneditor.php

OpenStudy (anonymous):

haha... found it... http://openstudy.com/study#/groups/latex practicing! :)

OpenStudy (anonymous):

Nice. Cool.

OpenStudy (anonymous):

Thanks, it actually makes sense!